# Boosting Generic Semi-Supervised Medical Image Segmentation via Diverse Teaching and Label Propagation

### v1

@@ IV-D2 Effectiveness of Diverse teaching
To generate reliable pseudo labels and and realize diverse consistency learning, we employ the entropy-based teacher ensemble to acquire an ensembled prediction. As presented in Tab. XI, when compared to using only one of the two teacher predictions, the ensembled diverse teaching can notably boost the model performance. Particularly on the Synapse dataset, this approach results in an approximately 2 increase in the Dice score.
TABLE XI: Effectiveness of Diverse teaching on the LA dataset with 5% labeled data and Synapse dataset using 20 % labeled data.
@@ IV-E2 Ablation Study of mask ratio
We further conduct ablation experiments on the Synapse dataset with 20% labeled data to analyze the parameter sensitivity of mask ratio $r$ in eq. 14. As shown in Tab. XII, the value of mask ratio $r$ that is either too high or too low may result in sub-optimal Dice performance. A relatively high mask ratio has the potential to obscure the majority of the foreground organs, thereby misleading the model during the training process. On the other hand, a low value of the mask ratio $r$ gives rise to challenges in learning the domain-invariant features to reduce the distribution shifts and capturing the local context within the image.
TABLE XII: Ablation Study of mask ratio on the Synapse dataset with 20% labeled data.
@@ V Conclusion
This paper introduces a novel unified approach, called Diverse Teaching and Label Propagation Network (DTLP-Net), for addressing the annotation-efficient medical image segmentation tasks, including classical SSMIS, UMDA, and Semi-MDA. Our DTLP-Net incorporates diverse teaching strategies to generate reliable pseudo-labels for the student model. Subsequently, it realizes global-local consistency learning through inter-sample and intra-sample data augmentation, guided by these pseudo-labels. In addition, masked reconstruction on the feature level and Knowledge distillation from the soft prediction is further utilized to alleviate the hard pseudo labels generated by the dual teachers. Ultimately, to fully exploit the potential of unlabeled data, a label propagation approach is put forward. This approach aims to learn pairwise similarities at the voxel-level for correlation consistency learning, thereby enhancing the model’s generalization ability. Extensive experiments carried out on five benchmark datasets have verified the effectiveness of the proposed methodology, indicating the potential of our framework to tackle more challenging SSL scenarios.

### v3

@@ IV-D2 Effectiveness of Diverse teaching
To generate reliable pseudo labels and and realize diverse consistency learning, we employ the entropy-based teacher ensemble to acquire an ensembled prediction. As presented in Tab. X, when compared to using only one of the two teacher predictions, the ensembled diverse teaching can notably boost the model performance. Particularly on the Synapse dataset, this approach results in an approximately 2% increase in the Dice score.
TABLE X: Effectiveness of Diverse teaching on the LA dataset with 5% labeled data and Synapse dataset using 20 % labeled data.
@@ IV-E2 Ablation Study of mask ratio
We further conduct ablation experiments on the Synapse dataset with 20% labeled data to analyze the parameter sensitivity of mask ratio $r$ in eq. 10. As shown in Tab. XI, the value of mask ratio $r$ that is either too high or too low may result in sub-optimal Dice performance. A relatively high mask ratio has the potential to obscure the majority of the foreground organs, thereby misleading the model during the training process. On the other hand, a low value of the mask ratio $r$ gives rise to challenges in learning the domain-invariant features to reduce the distribution shifts and capturing the local context within the image.
TABLE XI: Ablation Study of mask ratio on the Synapse dataset with 20% labeled data.
@@ V Conclusion
This paper introduces a novel unified approach, called Diverse Teaching and Label Propagation Network (DTLP-Net), for addressing the annotation-efficient medical image segmentation tasks, including classical SSMIS, UMDA, and Semi-MDA. Our DTLP-Net incorporates diverse teaching strategies to generate reliable pseudo-labels for the student model. Subsequently, it realizes global-local consistency learning through inter-sample and intra-sample data augmentation, guided by these pseudo-labels. In addition, masked reconstruction on the feature level and knowledge distillation from the soft prediction is further utilized to alleviate the hard pseudo labels generated by the dual teachers. Ultimately, to fully exploit the structure of the data, a label propagation approach is put forward. This approach aims to learn pairwise similarities at the voxel-level for correlation consistency learning, thereby enhancing the model’s generalization ability. Extensive experiments carried out on five benchmark datasets have verified the effectiveness of the proposed methodology, indicating the potential of our framework to tackle more challenging SSL scenarios.